OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 14 −20 36 −3 −11

OpenStudy (anonymous):

well if f(x) = 14 then f(x) = -2x +8 14 = -2x + 8 14 - 8 = -2x 6 = -2x 6/-2 = x -3 = x So, x = -3

OpenStudy (anonymous):

we put 14 instead of f(x):\[14=-2x+8 =>-2x=6=>x=-3\] hope it helps you.